# A Safe Preference Learning Approach for Personalization with Applications to Autonomous Vehicles

### 22 Jan 2024 v2

@@ IV-B Computational Approach
Thanks to the prevalence and success of gradient-based methods and back-propagation in machine learning, many temporal logic learning algorithms using gradients have been proposed [38]. To be able to compute the gradient, we need a differentiable loss function. In the WSTL robustness definition, we replace max and min functions with their soft differentiable versions softmin/softmax as in [38]. We also replace the indicator function with the logistic function with a shift. The shift helps for avoiding equality of robustness values in preference pairs. Overall, we propose the following surrogate loss
@@
(b) Pedestrian Scenario: Vehicle approaching to a pedestrian crosswalk, while a pedestrian is crossing. Vehicle can come to a complete stop or slow down sufficiently to allow the pedestrian.
@@
Implementation details: Inspired from [38], we construct a computation graph for robustness of WSTL formulas from syntax trees. This computation graph takes a signal as input and returns the WSTL robustness value of that signal at all times as output. We use PyTorch along with Adam [39] optimizer. Several strategies are investigated to mitigate convergence issues of the gradient-based method: (i) decreasing the softness coefficient $\beta$ of softmin/max, possibly compromising the soundness guarantee, (ii) decreasing the steepness of the logistic function, i.e., decreasing $M$, but this makes the surrogate $\mathcal{L}$ less similar to the objective in Problem 2, (iii) initializing the iteration from multiple random points to overcome bad local minima.
@@
Implementation details: We want the weight valuations to meet the following criterion: the absolute difference in robustness between signals within a pair should exceed $5\%$ of the range between the maximum and minimum robustness values among all signals. While this condition is not required for the random sampling approach alone, it can be useful for two downstream tasks: (i) when using the best performing of these weights as initialization of gradient-based approaches [^4] We tried this combination in our experiments, however, the performance improvement was not significant. Therefore, due to space constraints, we do not report these results further., this separation helps start the iterations at a part of the weight space where logistic function well-approximates the indicator function; (ii) when using the learned formula in controller synthesis, weights that well-separates the preferences lead to controllers that more robustly reflect the preferences.
@@ V Experiments
In this section, we provide a comparison of solution approaches with baseline methods, along with demonstrating the need for a safety-guaranteed preference learning framework. We also showcase the framework’s performance in capturing personal preferences of different participants in a human subject study. For these purposes, we use two different driving scenarios.
Driving Scenarios: We use STL to specify traffic rules in driving scenarios. The first scenario is a simple intersection with a stop sign, a screenshot is shown in Figure 1(a). The vehicle must stop before the stop sign, but there is some flexibility in the approach and final position. The traffic rule can be expressed as follows: $\phi^{stop}=\Diamond\square(x-x_{stop}\geq 0\wedge v=0)\wedge\square(v\geq 0)$ where $x$ and $v$ are the position and speed signals of a vehicle, respectively, and $x_{stop}$ is the stop sign position. Note that $\rho(s,\phi^{stop})\leq 0$ for any signal due to equality condition. We substitute $v=0$ with an indicator variable as discussed in Remark 1. We construct PWSTL formula $\phi^{stop}_{\mathcal{W}}$ with a weight parameter set $\mathcal{W}$ that contains all weights in the formula. In the second scenario, we observe an ego vehicle approaching to a pedestrian while she is crossing the road, as illustrated in Figure 1(b). The traffic regulation in this case is expressed in STL form as $\phi^{pedes}=\square[\big{(}p\wedge(x-x_{cross}\leq 0)\big{)}\implies\big{(}x-x_{cross}\leq 0\ \mathcal{U}\ \lnot p\big{)}\wedge(v\leq v_{lim})]$ where $x$, $v$ represent position and velocity signals, respectively. Boolean signal $p$ indicates the presence of a pedestrian, and $v_{lim}$ and $x_{cross}$ are constants denoting the speed limit and the crosswalk position, respectively.
Human Subject Studies: Studies are completed under IRB Study No. HUM00221976. For each scenario, we collaborate with eight participants with $75-25\%$ male-female ratio from the $25-35$ age group. We simulate hundred trajectories that satisfy the temporal logic formula per scenario. We compose fifty pairs such that the Euclidean distance between each pair is greater than a threshold. This threshold value is determined manually as the point at which the difference between signals becomes difficult to discern. These pairs are shown to participants who then choose their preferred behavior. As human decisions can vary in consistency, for the first scenario, we repeat the same question set twice to get a measure of the participant’s decisiveness. The consistency of the answers of the participants is reported in Figure 2(a).
@@
Figure 3: Human subject study results for the two scenarios for all users. “STL” denotes the traditional (unweighted) robustness when it is used directly,“RS” denotes our method with random sampling, “GB” denotes our method with gradient-based optimization, “BT” denotes SGD with Bradley-Terry model, and “SVM” represents SVM classification.
@@ V-A Baseline Methods
One well-known approach to pairwise preference learning problem is to recast it as a supervised learning problem [42]. Let $\psi(s)$ be the feature vector of item $s$. We construct $\psi(\cdot)$ by dividing the Fourier transform of $s$ into five frequency bins and adding the traditional robustness metric as the final feature. To set up the supervised learning problem, for a given preference pair $(s_{i}^{+},s_{i}^{-})$, we construct a new feature vector as the difference of feature vectors as $\psi(s_{i}^{+})-\psi(s_{i}^{-})$. All signal pairs in $\mathcal{P}$ belong to Class 00. We generate the data for Class $1$ by reversing the signal order and defining the feature vector $\psi(s_{i}^{-})-\psi(s_{i}^{+})$. This process gives us binary labels for all comparison pairs and their reverse orders. Then, we use Support Vector Machines (SVM) to learn a binary classifier. In the SVM, we use radial basis functions as kernels. For a test pair $(s_{1},s_{2})$, if $\psi(s_{1})-\psi(s_{2})$ is classified in Class 00, we say $s_{1}$ is preferred over $s_{2}$; and we say $s_{2}$ is preferred over $s_{1}$ otherwise.
@@
In particular, we use stochastic gradient decent (SGD) for solving this problem. Finally, for a test pair $(s_{1},s_{2})$, if $e^{<v^{*},\psi(s_{1})>}>e^{<v^{*},\psi(s_{2})>}$, we say $s_{1}$ is preferred over $s_{2}$; and we say $s_{2}$ is preferred over $s_{1}$ otherwise.
@@ V-B Comparison of Solution Approaches
In this section, we compare the performance of the proposed solution approaches with the baseline preference learning methods listed in Section V-A. We use the percentage of training (test) pairs that a model accurately predicts as the metric for comparison.
For each participant, we randomly split the preference set by $70\%-30\%$ as train-test data ten times, i.e., $35$ pairs for the training set and $15$ for the test set. For each train-test split, we compute the accuracy of train-test datasets with respect to traditional STL, and compare two proposed computational approaches with two baseline methods. The first method solves Problem 2 using $\phi^{stop}$ and $\phi^{pedes}$ for respective scenarios, via random sampling with a threshold condition, where we sample $1000$ weight valuations per split. For the stop sign scenario, the second method solves Problem 2 with gradient-based optimization over the loss function $\mathcal{L}$, initialized eleven times, ten from random weight valuations and one from the traditional STL valuation. We report the best training/test accuracy pair among these 11 as the result. The learning rate is $10^{-5}$, $\epsilon=0.01$, and $\theta=0.01$. The softness coefficient for softmax/min is $\beta=10^{10}$. We terminate the optimization when the cost value difference drops below $10^{-6}$. We divide the training set into batches of five pairs. Batch selection is random at each iteration. The third method is the SVM classification baseline. For the last method, we solve Equation (4) via SGD with the learning rate of $0.1$.
Results are shown in Figures 2(a) and 2(b). The average performance of all methods for all users and splits is shown in Table I. It can be seen that random sampling gives competitive results with the baseline methods in the stop sign scenario, and outperforms all methods in the pedestrian. Although gradient-based method gives equally good results as random sampling for the stop sign scenario, it is much slower. Indeed, it was too slow to converge for the pedestrian scenario, due to formula complexity, and is skipped. We conclude that simple random sampling effectively identifies promising weight valuations that improve the traditional STL accuracy, and gives comparable results to other methods.
@@
Finally, when we look at Figure 2(a), we see that with decreasing consistency, generalizability of all methods decreases, i.e., they perform poorly on test data.
Now we turn our attention to the safety of different approaches. Ideally, when presented with a pair of signals where one is violating the traffic rules and the other is satisfying, an approach should give preference to the satisfying one. Our method satisfies this nice property by construction. To test how the baselines do in this case, we simulate a hundred violating pairs for the intersection with stop sign scenario, and pair them with satisfying signals. Now, we have fifty satisfying-satisfying signal pairs that we use in human subject studies and a hundred satisfying-violating signal pairs. We create two different training sets: (i) one with fifty satisfying-satisfying pairs, and (ii) one with fifty satisfying-violating pairs in addition to satisfying-satisfying pairs. Test sets are hundred satisfying-violating pairs, and fifty satisfying-violating pairs, respectively. Table II shows the safety performance of two baseline methods, and random sampling for the stop sign scenario and for one participant. As we can see, baseline methods trained with only satisfying signals perform poorly when encountered with violating signals. However, it is not always feasible to generate real-life behaviors that violate a rule for safety-critical scenarios. When training baseline methods, we rely on simulators to generate violating signals, which may not be realistic. When we look at the results with training set (ii), the test performance of both baseline methods increases, and SVM reaches $100\%$ accuracy. However, none of the baseline methods ensure the safety for arbitrary safe-unsafe pairs.
@@ VI Conclusion, Limitations, and Future Work
We note that neither random-sampling nor gradient-based method guarantee finding an optimal value. We also observe the gradient-based method to have difficulties in convergence for certain formulas. It would be interesting to study different smooth robustness metrics to see if they can mitigate this issue. While preference data in our experiments appears to be on a smaller scale, expecting humans to select preferences for hundreds of signal pairs all at once is impractical. Our experience shows that even dealing with fifty pairs could be overwhelming. To this end, our upcoming focus is on an active learning scheme that maximizes inference using minimum amount of question pairs. Additionally, we aim to integrate the final WSTL formula into a downstream control synthesis algorithm as in [13] to demonstrate its use in control design and to run further validation studies.

### 12 Mar 2024 v3

@@ IV-B Computational Approach
Thanks to the prevalence and success of gradient-based methods and back-propagation in machine learning, many temporal logic learning algorithms using gradients have been proposed [39]. To be able to compute the gradient, we need a differentiable loss function. In the WSTL robustness definition, we replace max and min functions with their soft differentiable versions softmin/softmax as in [39]. We also replace the indicator function with the logistic function with a shift. The shift helps avoid equality of robustness values in preference pairs. Overall, we propose the following surrogate loss
@@
(b) Pedestrian Scenario: Vehicle approaching to a pedestrian crosswalk, while a pedestrian is crossing. The vehicle can come to a complete stop or slow down sufficiently to allow the pedestrian.
@@
Implementation details: Inspired by [39], we construct a computation graph for the robustness of WSTL formulas from syntax trees. This computation graph takes a signal as input and returns the WSTL robustness value of that signal at all times as output. We use PyTorch along with Adam [40] optimizer. Several strategies are investigated to mitigate convergence issues of the gradient-based method: (i) decreasing the softness coefficient $\beta$ of softmin/max, possibly compromising the soundness guarantee, (ii) decreasing the steepness of the logistic function, i.e., decreasing $M$, but this makes the surrogate $\mathcal{L}$ less similar to the objective in Problem 2, (iii) initializing the iteration from multiple random points to overcome bad local minima.
@@
Implementation details: We want weight valuations such that the absolute difference in robustness of signals within a pair should exceed $5\%$ of the range between the maximum and minimum robustness values among all signals. While this condition is not required for the random sampling approach alone, it can be useful for two downstream tasks: (i) when using the best performing of these weights as initialization of gradient-based approaches [^4] We tried this combination in our experiments, however, the performance improvement was not significant. Therefore, due to space constraints, we do not report these results further., this separation helps start the iterations at a part of the weight space where the logistic function well-approximates the indicator function; (ii) when using the learned formula in controller synthesis, weights that well-separates the preferences lead to controllers that more robustly reflect the preferences.
@@ V Experiments
In this section, we provide a comparison of solution approaches with baseline methods, along with demonstrating the need for a safety-guaranteed preference learning framework. [^5] The code and the data can be accessed from https://github.com/ruyakrgl/SPL-WSTL We also showcase the framework’s performance in capturing the personal preferences of different participants in a human subject study. For these purposes, we use two different driving scenarios.
Driving Scenarios: We use STL to specify traffic rules in driving scenarios. The first scenario is a simple intersection with a stop sign, a screenshot is shown in Figure 1(a). The vehicle must stop before the stop sign, but there is some flexibility in the approach and final position. The traffic rule can be expressed as follows: $\phi^{stop}=\Diamond\square(x-x_{stop}\geq 0\wedge v=0)\wedge\square(v\geq 0)$ where $x$ and $v$ are the position and speed signals of a vehicle, respectively, and $x_{stop}$ is the stop sign position. Note that $\rho(s,\phi^{stop})\leq 0$ for any signal due to equality condition. We substitute $v=0$ with an indicator variable as discussed in Remark 1. We construct the PWSTL formula $\phi^{stop}_{\mathcal{W}}$ with a weight parameter set $\mathcal{W}$ that contains all weights in the formula. In the second scenario, we observe an ego vehicle approaching a pedestrian while she is crossing the road, as illustrated in Figure 1(b). The traffic regulation, in this case, is expressed in STL form as $\phi^{pedes}=\square[\big{(}p\wedge(x-x_{cross}\leq 0)\big{)}\implies\big{(}x-x_{cross}\leq 0\ \mathcal{U}\ \lnot p\big{)}\wedge(v\leq v_{lim})]$ where $x$, $v$ represent position and velocity signals, respectively. Boolean signal $p$ indicates the presence of a pedestrian, and $v_{lim}$ and $x_{cross}$ are constants denoting the speed limit and the crosswalk position, respectively.
Human Subject Studies: Studies are completed under IRB Study No. HUM00221976. For each scenario, we collaborate with eight participants with a $75-25\%$ male-female ratio from the $25-35$ age group. We simulate a hundred trajectories that satisfy the temporal logic formula per scenario. We compose fifty pairs such that the Euclidean distance between each pair is greater than a threshold. This threshold value is determined manually as the point at which the difference between signals becomes difficult to discern. These pairs are shown to participants who then choose their preferred behavior. As human decisions can vary in consistency, for the first scenario, we repeat the same question set twice to get a measure of the participant’s decisiveness. The consistency levels of four participants are reported in Figure 2(a).
@@
Figure 3: Human subject study results for the two scenarios for four of the users. “STL” denotes the traditional (unweighted) robustness when it is used directly, “RS” denotes our method with random sampling, “GB” denotes our method with gradient-based optimization, “BT” denotes SGD with Bradley-Terry model, and “SVM” represents SVM classification.
@@ V-A Baseline Methods
One well-known approach to pairwise preference learning problem is to recast it as a supervised learning problem [43]. Let $\psi(s)$ be the feature vector of item $s$. We construct $\psi(\cdot)$ by dividing the Fourier transform of $s$ into five frequency bins and adding the traditional robustness metric as the final feature. To set up the supervised learning problem, for a given preference pair $(s_{i}^{+},s_{i}^{-})$, we construct a new feature vector as the difference of feature vectors as $\psi(s_{i}^{+})-\psi(s_{i}^{-})$. All signal pairs in $\mathcal{P}$ belong to Class 00. We generate the data for Class $1$ by reversing the signal order and defining the feature vector $\psi(s_{i}^{-})-\psi(s_{i}^{+})$. This process gives us binary labels for all comparison pairs and their reverse orders. Then, we use Support Vector Machines (SVM), with a radial basis function kernel, to learn a binary classifier. For a test pair $(s_{1},s_{2})$, if $\psi(s_{1})-\psi(s_{2})$ is classified in Class 00, we say $s_{1}$ is preferred over $s_{2}$; and we say $s_{2}$ is preferred over $s_{1}$ otherwise.
@@
In particular, we use stochastic gradient descent (SGD) for solving this problem. Finally, for a test pair $(s_{1},s_{2})$, if $e^{<v^{*},\psi(s_{1})>}>e^{<v^{*},\psi(s_{2})>}$, we say $s_{1}$ is preferred over $s_{2}$; and we say $s_{2}$ is preferred over $s_{1}$ otherwise.
@@ V-B Comparison of Solution Approaches
In this section, we compare the performance of the proposed solution approaches with the baseline methods listed in Section V-A. We use the percentage of training (test) pairs that a model accurately predicts as the metric for comparison.
For each participant, we use ten random $70\%-30\%$ splits of the preference set in to train-test data, i.e., $35$ pairs for the training set and $15$ for the test set. For each split, we compute the train-test accuracy with respect to traditional STL and compare two proposed approaches with two baseline methods. Our first method solves Problem 2 using $\phi^{stop}$ and $\phi^{pedes}$ for respective scenarios, via random sampling with a threshold condition, where we sample $1000$ weight valuations per split. For the stop sign scenario, our second method solves Problem 2 with gradient-based optimization over the loss function $\mathcal{L}$, initialized eleven times, ten from random weight valuations and one from the traditional STL valuation. We report the best training/test accuracy pair among these 11 as a result. The learning rate is $10^{-5}$, $\epsilon=0.01$, and $\theta=0.01$. The softness coefficient for softmax is $\beta=10^{10}$. We terminate the optimization when the cost value difference drops below $10^{-6}$. We divide the training set into batches of five pairs. Batch selection is random at each iteration. The third method is the SVM classification baseline. For the last method, we solve Equation (4) via SGD with the learning rate of $0.1$.
Some representative results are shown in Figures 2(a) and 2(b). The average performance of all methods for all users and splits is shown in Table I. Random sampling gives competitive results with the baseline methods in the stop sign scenario, and outperforms all methods in the pedestrian. Although the gradient-based method gives equally good results as random sampling for the stop sign scenario, it is much slower. Indeed, it was too slow to converge for the pedestrian scenario, due to formula complexity, and is skipped. We conclude that simple random sampling effectively identifies promising weight valuations that improve the traditional STL accuracy, and give comparable results to other methods.
@@
Finally, when we look at Figure 2(a), we see that with decreasing consistency, the generalizability of all methods decreases, i.e., they perform poorly on test data.
Now we turn our attention to the safety of different approaches. Ideally, when presented with a pair of signals where one is violating the traffic rules and the other is satisfying, an approach should give preference to the satisfying one. Our method satisfies this nice property by construction. To test how the baselines do in this case, we simulate a hundred violating pairs for the intersection with a stop sign scenario and pair them with satisfying signals. Now, we have fifty satisfying-satisfying signal pairs that we use in human subject studies and a hundred satisfying-violating signal pairs. We create two different training sets: (i) one with fifty satisfying-satisfying pairs, and (ii) one with fifty satisfying-violating pairs in addition to satisfying-satisfying pairs. Test sets are hundred satisfying-violating pairs, and fifty satisfying-violating pairs, respectively. Table II shows the safety performance of two baseline methods and random sampling for the stop sign scenario and one participant. As we can see, baseline methods trained with only satisfying signals perform poorly when encountered with violating signals. However, it is not always feasible to generate real-life behaviors that violate a rule for safety-critical scenarios. When training baseline methods, we rely on simulators to generate violating signals, which may not be realistic. When we look at the results with training set (ii), the test performance of both baseline methods increases, and SVM reaches $100\%$ accuracy. However, none of the baseline methods ensure the safety of arbitrary safe-unsafe pairs.
@@ VI Conclusion, Limitations, and Future Work
We note that neither random sampling nor gradient-based method guarantees finding an optimal value. We also observe the gradient-based method to have difficulties in convergence for certain formulas. It would be interesting to study different smooth robustness metrics to see if they can mitigate this issue. While preference data in our experiments appears to be on a smaller scale, expecting humans to select preferences for hundreds of signal pairs all at once is impractical. Our experience shows that even dealing with fifty pairs could be overwhelming. To this end, our upcoming focus is on an active learning scheme that maximizes inference using a minimum amount of question pairs. Additionally, we aim to integrate the final WSTL formula into a downstream control synthesis algorithm as in [13] to demonstrate its use in control design and to run further validation studies.